# Efficient Perplexity Bound and Ratio Matching in Discrete Diffusion Language Models

### v1

@@ 1 Introduction
Modeling data distributions is a fundamental task in machine learning. In the case of continuous data distributions, recent advancements in continuous diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2020b) have demonstrated impressive capabilities in generating data samples and performing density estimation (Song et al., 2021a; c; Haxholli & Lorenzi, 2023; Kingma et al., 2021a). Despite these achievements, the application of such models to categorical data distributions, like language, remains limited, as continuous diffusion models generally underperform compared to autoregressive models in these scenarios (Chen et al., 2023; Gulrajani & Hashimoto, 2024; Li et al., 2022; Dieleman et al., 2022; Strudel et al., 2022). To address this, recent research has focused on the development of discrete diffusion models (Austin et al., 2021; Campbell et al., 2022; Meng et al., 2022; Lou et al., 2024; Sahoo et al., 2024; Shi et al., 2024; Ou et al., 2025) which offer distinct advantages compared to autoregressive models, such as the ability to infill various parts of a sequence non-sequentially and have the potential to reduce computing time and expenses in generating lengthy sequences (Deschenaux & Gulcehre, 2024; Christopher et al., 2024).
@@
To illustrate our point, we design a new transition-rate matrix named roulette diffusion, and derive its matrix exponential. The roulette diffusion is an interpolation between the absorb and the uniform diffusion. In the forward process, a token can transition to any state until it hits the absorb state, that is, until it is masked. In return, the reverse process begins with a sequence of masked tokens, which are gradually unmasked, and where the unmasked tokens can be refined. Intuitively. this capability should be important for discrete diffusion models (Deschenaux & Gulcehre, 2024), and can be useful for downstream tasks, as shown in our spelling correction experiments. Moreover, during the reverse process, the scores/ratios of unmasked tokens have much larger magnitudes than those of masked tokens, posing a significant learning challenge for the network due to output scale variability. Employing the CEDD strategy mitigates this challenge as demonstrated experimentally.
@@ 2.1 Markov Chains Over Finite-State Spaces
One can generalize such processes into Continuous Time Markov Chains (CTMCs) where $t\in[0,\bar{T}]$, (Anderson, 2012). For simplicity, we make the choice $\bar{T}=1$. To construct a CTMC, one first chooses a transition-rate matrix ${\bm{Q}}_{t}$, which has the property that its non-diagonal elements are non-negative, and the elements in each of its columns add to zero (Suhov & Kelbert, 2008). Given an initial probability distribution ${\bm{s}}_{0}$, the equation below fully determines the evolution of the probability with respect to time:
@@ 2.3 Discrete Diffusion for Language Modeling - Token Level Transitions
$p(x_{t-\epsilon}={\bm{y}}\mid x_{t}={\bm{x}})=\begin{cases}1-\sum_{z^{i}\in\text{Vocab}\setminus x^{i}}{\bm{Q}}^{tok}_{t}(x^{i},z^{i})\frac{p_{t}({\bm{z}})}{p_{t}({\bm{x}})}\epsilon+O(\epsilon^{2}).&\text{if }{\bm{y}}={\bm{x}}\\ {\bm{Q}}^{tok}_{t}(x^{i},y^{i})\frac{p_{t}({\bm{y}})}{p_{t}({\bm{x}})}\epsilon+O(\epsilon^{2}).&\text{if }{\bm{y}}\neq{\bm{x}}\\ \end{cases}$ (8)
@@
Finally, the noise schedule $\sigma_{t}$ is typically loglinear $-\log(1-(1-\epsilon)t)$ or geometric $\sigma_{min}^{1-t}\cdot\sigma_{max}^{t}$.
@@ 3.1 Cross Entropy and KL Divergence Results
Define a CTMC with transition matrix ${{\bm{Q}}}_{t}$ that runs from time 00 to $1$. The true reverse process defines a probability evolution $p_{t}$ from $p_{1}$ to the data distribution $p_{0}$, while the learned reverse process induces the evolution $p^{\theta}_{t}$ from the reference distribution $p^{\theta}_{1}=p_{r}$ to the approximation of the data distribution $p^{\theta}_{0}$. In this setting, the following KL divergence bound holds
@@
$-\log p_{0}^{\theta}(x_{0})\leq\int_{0}^{1}\mathbb{E}_{x_{t}\sim p_{t|0}(\cdot|x_{0})}\sum_{y\neq x_{t}}{\bm{Q}}_{t}(x_{t},y)\ell\left(\frac{p_{t|0}(y|x_{0})}{p_{t|0}(x_{t}|x_{0})},s_{\theta}(x_{t},t)_{y}\right)dt$
$+D_{KL}(p_{1|0}(\cdot|x_{0})||p_{r}).$ (12)
@@
$-\int_{0}^{1}\mathbb{E}_{x_{t}\sim p_{t}}\sum_{y\neq x_{t}}{\bm{Q}}_{t}(y,x_{t})dt+H(p_{1},p_{r}),\text{ where }\bar{\ell}(a,b)=\left(b-a\log b\right).$ (14)
The second term $-\int_{0}^{1}\mathbb{E}_{x_{t}\sim p_{t}}\sum_{y\neq x_{t}}{\bm{Q}}_{t}(y,x_{t})dt$ and third one $H(p_{1},p_{r})\approx H(p_{r})$ can be analytically computed as shown in Section 3.3, Proposition 6. Finally, the first term can be rewritten as
@@ 3.2 Roulette Discrete Diffusion
Typically, matrices ${\bm{Q}}^{tok}$ are defined as ${\bm{Q}}^{tok}={\bm{P}}^{tok}-{\bm{I}}$, where ${\bm{P}}^{tok}$ is idempotent, since this implies that $({\bm{Q}}^{tok})^{2}=-{\bm{Q}}^{tok}$. This last property of ${\bm{Q}}^{tok}$ greatly simplifies the calculation of $e^{\sigma(t){\bm{Q}}^{tok}}$, as by using the Taylor series, $e^{\sigma(t){\bm{Q}}^{tok}}={\bm{I}}+{\bm{Q}}^{tok}(1-e^{-\sigma(t)})$. Usually, the following two matrices ${\bm{P}}^{tok}$ are chosen: a) ${\bm{P}}^{tok}_{uniform}(V\times V)$ where each entry is set to $\frac{1}{V}$, and b) ${\bm{P}}^{tok}_{absorb}(n\times n)$ in which the last row is full of ones while all other elements are 00, where $n=V+1$. While absorb diffusion often outperforms the uniform one in standard evaluations, the latter is more practical for some tasks like spelling correction, where refining tokens is crucial. We propose another transition-rate matrix whose exponential can be analytically calculated. To our knowledge this is the only matrix with such a property presented so far apart from the absorb and uniform ones. We refer to this new discrete diffusion process as roulette diffusion. The corresponding ${\bm{P}}^{tok}_{roulette}(n\times n)$ is a matrix, such that ${\bm{P}}^{tok}_{roulette}(i\neq n,j\neq n)=\frac{1}{V}(1-p_{m})$, ${\bm{P}}^{tok}_{roulette}(n,j\neq n)=p_{m}$, ${\bm{P}}^{tok}_{roulette}(i\neq n,n)=0$ and ${\bm{P}}^{tok}_{roulette}(n,n)=1$. We notice that for $p_{m}=1$, roulette diffusion coincides with absorb diffusion, while for $p_{m}=0$ it coincides with the uniform diffusion. Intuitively, a token can transit from a non-absorb state to a non-absorb state with probability $\frac{1}{V}(1-p_{m})$, until it hits the absorb state (with probability $p_{m}$) and then remains there. While this matrix is not idempotent, one can still calculate its exponential as stated in the following proposition (proved in Appendix A.3.3):
@@
The noise schedule used is the roulette-loglinear noise $-\frac{1}{p_{m}}\log(1-(1-\epsilon)t)$. In the reverse process, when a token is unmasked it can still be corrected with probability directly related to $p_{m}$ as shown in Appendix A.3.4. A generalization for time-evolving $p_{m}$ is given in Appendix A.3.5, and the corresponding ${\bm{Q}}^{tok}_{eroulette}(t)$ is named eroulette. Therein (Proposition 7), it is shown that the exponential matrix can be calculated as in the previous proposition, by substituting $p_{m}$ with $p_{m}(t)$.
@@ 3.3 Efficient Estimation of $J_{2}$
[…]
@@ 3.4 Modeling Ratios via CEDD
In the uniform case, models using the direct reparametrization in (18) underperform SEDD in terms of perplexity. Indeed, the model is too confident in its predictions when $t\rightarrow 0$ as it does not benefit form the neural network regularization due to the incorporation of the true conditional ratios. Thus, inside the ratios $\frac{p_{t|0}({\bm{y}}^{i}|h)}{p_{t|0}({{\bm{x}}_{t}}^{i}|h)}$ we rescale $\sigma^{\theta}_{t}<0.0015$, by setting $\sigma^{\theta}_{t}=0.0015$. We also perform rescaling in the case of roulette diffusion dynamics, as when time is close to 00 most tokens are unmasked. The same strategy is applied, as before, only to ratios corresponding to unmasked token, by rescaling $\sigma^{\theta}_{t}$ when $\sigma^{\theta}_{t}<0.5$ as follows: $\sigma^{scaled,\theta}_{t}=\log(1.1\sigma^{\theta}_{t}+1.1)$. Naturally, for the sake of rigor, these are also the models we employ to generate samples, whose quality is measured in terms of generative perplexity. Further, in order to evade metric hacking, we take care to not modify the $\sigma_{t}$ in the metrics $J_{1}$ and $J_{2}$ used to evaluate the models. Additional details and motivations for such rescalings are provided in the last paragraph of Appendix B.3.3. When $w(t)=1$ we refer to the method as CEDD, while when $w(t)=\log{(e+\frac{0.3}{t})}$ we use CEDD*.
@@ 4 Experiments
We now empirically validate the approaches and theoretical contributions presented in the previous section. In Subsection 4.1, we compare the generative perplexities of models trained on OpenWebText (Gokaslan & Cohen, 2019) with SEDD and SEDD scaled (SEDDs, see Appendix B.3.2 for details) versus those trained with CEDD and CEDD*. We keep all other variables unchanged for a fair comparison, finding that CEDD outperforms SEDD in all cases. The tests are conducted for the absorb, uniform and roulette diffusion dynamics. In Subsection 4.2, we evaluate the perplexity of the models, by calculating the upper bound on 5 different datasets, namely: 1BW, LAMBADA, PTB, Wikitext2 and Wikitext103 (Chelba et al., 2013; Paperno et al., 2016; Marcus et al., 1993; Merity et al., 2016). For the sake of reproducing the results of previous work we use $J_{1}$ as a metric. Finally, having computed the results using $J_{1}$, in Subsection 4.3 we also re-evaluate the models via $J_{2}$. Our findings suggest that the bound provided by $J_{2}$ is slightly tighter. In the last subsection we compare SEDD and CEDD* on a spell-checking task.
@@ 4.1 Generative Perplexity Comparisons
Initially, we perform a grid search to find the best $p_{m}$ for the roulette case. Out of the 4 values, 0.95, 0.65, 0.35, 0.050.950.650.350.050.95,\ 0.65,\ 0.35,\ 0.050.95 , 0.65 , 0.35 , 0.05, we found $p_{m}=0.95$ performs best (Appendix B.6, Table 5), but the optimum is likely reached when $p_{m}=1$. However, $p_{m}=0.95$, enables equipping unmasking dynamics with the correction mechanism, which can be useful for some tasks like spelling correction. Therefore, in what follows, when the roulette case is concerned, it is implied that $p_{m}=0.95$. In Table 1, we provide the results when the sequence length is 128 and the number of reverse steps is 1024, when using the analytic sampler. Generated samples can be found in Appendix C.1.
@@ 5 Related Work and Future Outlook
The continuous diffusion approach has demonstrated excellent performance in modeling continuous data distributions (Song et al., 2020a; 2021c; Kingma et al., 2021b; Nichol & Dhariwal, 2021; Saharia et al., 2022; Ramesh et al., 2022), leading to numerous efforts to adapt it for language modeling tasks (Chen et al., 2023; Gulrajani & Hashimoto, 2024; Li et al., 2022; Dieleman et al., 2022; Strudel et al., 2022; Gong et al., 2022; Mahabadi et al., 2023). Although initial results were not competitive, recent advancements have reduced the performance gap with autoregressive models. Discrete diffusion models offer an alternative approach for modeling categorical distributions like language data. Originally proposed by Hoogeboom et al. (2021); Austin et al. (2021), the framework was extended to continuous time by Campbell et al. (2022). Both strands of work employ a combination of the variational lower bound and cross entropy loss, the latter being central to our training approach and corresponding to the strategy employed by Dieleman et al. (2022) in the continuous case. The cross-entropy loss is also derived inSahoo et al. (2024); Shi et al. (2024); Ou et al. (2025), but only for the absorb transition-rate matrix. In addition, the cross entropy loss is similar to the loss employed in Sun et al. (2023), however in their case, one conditions on the current sequence ${\bm{x}}_{t}^{-i}$ without the token at position $i$, and attempts to maximize the likelihood of ${\bm{x}}^{i}_{t}$. In contrast, Lou et al. (2024) proposed modeling ratios of probabilities (Meng et al., 2022) directly using score entropy.

### v2

@@ 1 Introduction
Modeling data distributions is a fundamental task in machine learning. In the case of continuous data distributions, recent advancements in continuous diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2020b) have demonstrated impressive capabilities in generating data samples and performing density estimation (Song et al., 2021a; c; Haxholli and Lorenzi, 2023; Kingma et al., 2021b). Despite these achievements, the application of such models to categorical data distributions, like language, remains limited, as continuous diffusion models generally underperform compared to autoregressive models in these scenarios (Chen et al., 2023; Gulrajani and Hashimoto, 2024; Li et al., 2022; Dieleman et al., 2022; Strudel et al., 2022). To address this, recent research has focused on the development of discrete diffusion models (Austin et al., 2021; Campbell et al., 2022; Meng et al., 2022; Lou et al., 2024; Sahoo et al., 2024; Shi et al., 2024; Ou et al., 2025) which offer distinct advantages compared to autoregressive models, such as the ability to infill various parts of a sequence non-sequentially and have the potential to reduce computing time and expenses in generating lengthy sequences (Deschenaux and Gulcehre, 2024; Christopher et al., 2024).
@@
To illustrate our point, we design a new transition-rate matrix named roulette diffusion, and derive its matrix exponential. The roulette diffusion is an interpolation between the absorb and the uniform diffusion. In the forward process, a token can transition to any state until it hits the absorb state, that is, until it is masked. In return, the reverse process begins with a sequence of masked tokens, which are gradually unmasked, and where the unmasked tokens can be refined. Intuitively. this capability should be important for discrete diffusion models (Deschenaux and Gulcehre, 2024), and can be useful for downstream tasks, as shown in our spelling correction experiments. Moreover, during the reverse process, the scores/ratios of unmasked tokens have much larger magnitudes than those of masked tokens, posing a significant learning challenge for the network due to output scale variability. Employing the CEDD strategy mitigates this challenge as demonstrated experimentally.
@@ 2.1 Markov Chains Over Finite-State Spaces
One can generalize such processes into Continuous Time Markov Chains (CTMCs) where $t\in[0,\bar{T}]$, (Anderson, 2012). For simplicity, we make the choice $\bar{T}=1$. To construct a CTMC, one first chooses a transition-rate matrix ${\bm{Q}}_{t}$, which has the property that its non-diagonal elements are non-negative, and the elements in each of its columns add to zero (Suhov and Kelbert, 2008). Given an initial probability distribution ${\bm{s}}_{0}$, the equation below fully determines the evolution of the probability with respect to time:
@@ 2.3 Discrete Diffusion for Language Modeling - Token Level Transitions
[…]
@@
Finally, the noise schedule $\sigma_{t}$ is typically loglinear $\log(1-(1-\epsilon)t)$ or geometric $\sigma_{min}^{1-t}\cdot\sigma_{max}^{t}$.
@@ 3.1 Cross Entropy and KL Divergence Results
Define a CTMC with transition matrix ${{\bm{Q}}}_{t}$ that runs from time 0 to $1$. The true reverse process defines a probability evolution $p_{t}$ from $p_{1}$ to the data distribution $p_{0}$, while the learned reverse process induces the evolution $p^{\theta}_{t}$ from the reference distribution $p^{\theta}_{1}=p_{r}$ to the approximation of the data distribution $p^{\theta}_{0}$. In this setting, the following KL divergence bound holds
@@
$\log p_{0}^{\theta}(x_{0})\leq\int_{0}^{1}\mathbb{E}_{x_{t}\sim p_{t|0}(\cdot|x_{0})}\sum_{y\neq x_{t}}{\bm{Q}}_{t}(x_{t},y)\ell\left(\frac{p_{t|0}(y|x_{0})}{p_{t|0}(x_{t}|x_{0})},s_{\theta}(x_{t},t)_{y}\right)dt$
$D_{KL}(p_{1|0}(\cdot|x_{0})||p_{r}).$ (12)
@@
[…]
The second term $\int_{0}^{1}\mathbb{E}_{x_{t}\sim p_{t}}\sum_{y\neq x_{t}}{\bm{Q}}_{t}(y,x_{t})dt$ and third one $H(p_{1},p_{r})\approx H(p_{r})$ can be analytically computed as shown in Section 3.3, Proposition 6. Finally, the first term can be rewritten as
@@ 3.2 Roulette Discrete Diffusion
Typically, matrices ${\bm{Q}}^{tok}$ are defined as ${\bm{Q}}^{tok}={\bm{P}}^{tok}-{\bm{I}}$, where ${\bm{P}}^{tok}$ is idempotent, since this implies that $({\bm{Q}}^{tok})^{2}=-{\bm{Q}}^{tok}$. This last property of ${\bm{Q}}^{tok}$ greatly simplifies the calculation of $e^{\sigma(t){\bm{Q}}^{tok}}$, as by using the Taylor series, $e^{\sigma(t){\bm{Q}}^{tok}}={\bm{I}}+{\bm{Q}}^{tok}(1-e^{-\sigma(t)})$. Usually, the following two matrices ${\bm{P}}^{tok}$ are chosen: a) ${\bm{P}}^{tok}_{uniform}(V\times V)$ where each entry is set to $\frac{1}{V}$, and b) ${\bm{P}}^{tok}_{absorb}(n\times n)$ in which the last row is full of ones while all other elements are 0, where $n=V+1$. While absorb diffusion often outperforms the uniform one in standard evaluations, the latter is more practical for some tasks like spelling correction, where refining tokens is crucial. We propose another transition-rate matrix whose exponential can be analytically calculated. To our knowledge this is the only matrix with such a property presented so far apart from the absorb and uniform ones. We refer to this new discrete diffusion process as roulette diffusion. The corresponding ${\bm{P}}^{tok}_{roulette}(n\times n)$ is a matrix, such that ${\bm{P}}^{tok}_{roulette}(i\neq n,j\neq n)=\frac{1}{V}(1-p_{m})$, ${\bm{P}}^{tok}_{roulette}(n,j\neq n)=p_{m}$, ${\bm{P}}^{tok}_{roulette}(i\neq n,n)=0$ and ${\bm{P}}^{tok}_{roulette}(n,n)=1$. We notice that for $p_{m}=1$, roulette diffusion coincides with absorb diffusion, while for $p_{m}=0$ it coincides with the uniform diffusion. Intuitively, a token can transit from a non-absorb state to a non-absorb state with probability $\frac{1}{V}(1-p_{m})$, until it hits the absorb state (with probability $p_{m}$) and then remains there. While this matrix is not idempotent, one can still calculate its exponential as stated in the following proposition (proved in Appendix A.3.3):
@@
The noise schedule used is the roulette-loglinear noise $\frac{1}{p_{m}}\log(1-(1-\epsilon)t)$. In the reverse process, when a token is unmasked it can still be corrected with probability directly related to $p_{m}$ as shown in Appendix A.3.4. A generalization for time-evolving $p_{m}$ is given in Appendix A.3.5, and the corresponding ${\bm{Q}}^{tok}_{eroulette}(t)$ is named eroulette. Therein (Proposition 7), it is shown that the exponential matrix can be calculated as in the previous proposition, by substituting $p_{m}$ with $p_{m}(t)$.
@@ 3.3 Efficient Estimation of $J_{2}$
In the case of the roulette diffusion with roulette-loglinear noise, $H(p_{r})=0$ and $\int_{0}^{1}\mathbb{E}_{{\bm{x}}_{t}\sim p_{t}}\sum_{{\bm{y}}\neq{\bm{x}}_{t}}{\bm{Q}}_{t}({\bm{y}},{\bm{x}}_{t})dt=\left(1-\frac{1-p_{m}}{n-1}\right)\frac{L}{p_{m}}(\epsilon-1)$. For the absorb diffusion, we have $H(p_{r})=0$ and $\int_{0}^{1}\mathbb{E}_{{\bm{x}}_{t}\sim p_{t}}\sum_{{\bm{y}}\neq{\bm{x}}_{t}}{\bm{Q}}_{t}({\bm{y}},{\bm{x}}_{t})dt=L(\epsilon-1)$. Finally, in the case of uniform diffusion, $H(p_{r})=L\log(V)$ and $\int_{0}^{1}\mathbb{E}_{{\bm{x}}_{t}\sim p_{t}}\sum_{{\bm{y}}\neq{\bm{x}}_{t}}{\bm{Q}}_{t}({\bm{y}},{\bm{x}}_{t})dt=-\left(1-\frac{1}{V}\right)L\int_{0}^{1}\sigma^{{}^{\prime}}(t)dt$.
@@ 3.4 Modeling Ratios via CEDD
In the uniform case, models using the direct reparametrization in (18) underperform SEDD in terms of perplexity. Indeed, the model is too confident in its predictions when $t\rightarrow 0$ as it does not benefit form the neural network regularization due to the incorporation of the true conditional ratios. Thus, inside the ratios $\frac{p_{t|0}({\bm{y}}^{i}|h)}{p_{t|0}({{\bm{x}}_{t}}^{i}|h)}$ we rescale $\sigma^{\theta}_{t}<0.0015$, by setting $\sigma^{\theta}_{t}=0.0015$. We also perform rescaling in the case of roulette diffusion dynamics, as when time is close to 0 most tokens are unmasked. The same strategy is applied, as before, only to ratios corresponding to unmasked token, by rescaling $\sigma^{\theta}_{t}$ when $\sigma^{\theta}_{t}<0.5$ as follows: $\sigma^{scaled,\theta}_{t}=\log(1.1\sigma^{\theta}_{t}+1.1)$. Naturally, for the sake of rigor, these are also the models we employ to generate samples, whose quality is measured in terms of generative perplexity. Further, in order to evade metric hacking, we take care to not modify the $\sigma_{t}$ in the metrics $J_{1}$ and $J_{2}$ used to evaluate the models. Additional details and motivations for such rescalings are provided in the last paragraph of Appendix B.3.3. When $w(t)=1$ we refer to the method as CEDD, while when $w(t)=\log{(e+\frac{0.3}{t})}$ we use CEDD*.
@@ 4 Experiments
We now empirically validate the approaches and theoretical contributions presented in the previous section. In Subsection 4.1, we compare the generative perplexities of models trained on OpenWebText (Gokaslan and Cohen, 2019) with SEDD and SEDD scaled (SEDDs, see Appendix B.3.2 for details) versus those trained with CEDD and CEDD*. We keep all other variables unchanged for a fair comparison, finding that CEDD outperforms SEDD in all cases. The tests are conducted for the absorb, uniform and roulette diffusion dynamics. In Subsection 4.2, we evaluate the perplexity of the models, by calculating the upper bound on 5 different datasets, namely: 1BW, LAMBADA, PTB, Wikitext2 and Wikitext103 (Chelba et al., 2013; Paperno et al., 2016; Marcus et al., 1993; Merity et al., 2016). For the sake of reproducing the results of previous work we use $J_{1}$ as a metric. Finally, having computed the results using $J_{1}$, in Subsection 4.3 we also re-evaluate the models via $J_{2}$. Our findings suggest that the bound provided by $J_{2}$ is slightly tighter. In the last subsection we compare SEDD and CEDD* on a spell-checking task.
@@ 4.1 Generative Perplexity Comparisons
Initially, we perform a grid search to find the best $p_{m}$ for the roulette case. Out of the 4 values, 0.95, 0.65, 0.35, 0.050.95,\ 0.65,\ 0.35,\ 0.05, we found $p_{m}=0.95$ performs best (Appendix B.6, Table 5), but the optimum is likely reached when $p_{m}=1$. However, $p_{m}=0.95$, enables equipping unmasking dynamics with the correction mechanism, which can be useful for some tasks like spelling correction. Therefore, in what follows, when the roulette case is concerned, it is implied that $p_{m}=0.95$. In Table 1, we provide the results when the sequence length is 128 and the number of reverse steps is 1024, when using the analytic sampler. Generated samples can be found in Appendix C.1.
@@ 5 Related Work and Future Outlook
The continuous diffusion approach has demonstrated excellent performance in modeling continuous data distributions (Song et al., 2020a; 2021c; Kingma et al., 2021a; Nichol and Dhariwal, 2021; Saharia et al., 2022; Ramesh et al., 2022), leading to numerous efforts to adapt it for language modeling tasks (Chen et al., 2023; Gulrajani and Hashimoto, 2024; Li et al., 2022; Dieleman et al., 2022; Strudel et al., 2022; Gong et al., 2022; Mahabadi et al., 2023). Although initial results were not competitive, recent advancements have reduced the performance gap with autoregressive models. Discrete diffusion models offer an alternative approach for modeling categorical distributions like language data. Originally proposed by Hoogeboom et al. (2021); Austin et al. (2021), the framework was extended to continuous time by Campbell et al. (2022). Both strands of work employ a combination of the variational lower bound and cross entropy loss, the latter being central to our training approach and corresponding to the strategy employed by Dieleman et al. (2022) in the continuous case. The cross-entropy loss is also derived inSahoo et al. (2024); Shi et al. (2024); Ou et al. (2025), but only for the absorb transition-rate matrix. In addition, the cross entropy loss is similar to the loss employed in Sun et al. (2023), however in their case, one conditions on the current sequence ${\bm{x}}_{t}^{-i}$ without the token at position $i$, and attempts to maximize the likelihood of ${\bm{x}}^{i}_{t}$. In contrast, Lou et al. (2024) proposed modeling ratios of probabilities (Meng et al., 2022) directly using score entropy.